OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? x − y = 6 4x − 3y = 12 Add the two equations together Multiply the first equation by −4 Multiply the first equation by 3 Multiply the first equation by −3

OpenStudy (anonymous):

Fourth option multiply the first equation by -3 because x − y = 6 when we multiply it becomes -3(x-y=6) -3x+3y=-18 When we add them y can be eliminated -3x+3y=-18 4x − 3y = 12 ------------ x=-6 I hope I'm clear :)